a rectangular field is twice as long as it is wide. If the perimeter of th field is 72 yards, what are its dimensions
Answer by checkley77(12844)

L=2W
2L+2W=72
2*2W+2W=72
4W+2W=72
6W=72
W=72/6
W=12 ANS. FOR THE WIDTH.
L=2*12=24 ANS FOR THE LENGTH.
PROOF:
2*24+2*12=72
48+24=72
72=72